# VLBI astrometry of radio stars to link radio and optical celestial reference frames III: 11 radio stars

### v1

@@ 4.1 Gains of MultiView
The performance comparison between the sMV and PR can be conducted using the RMS-based JMFIT error or the fractional flux recovery (FFR) quantity, which is defined as the ratio between the peak brightness in the astrometric map and the self-calibrated map of the same source (2017AJ....153..105R). The radio stars are too weak for self-calibration, so here we use the peak flux density instead.
@@
Fig. 3 shows the ratio of sMV to PR versus the angular separation between the target and the primary calibrator. Similar to the JMFIT error, the sMV reduces flux loss in most cases (sMV/PR > 1, ), and the poor cases are mainly attributed to DM UMa and RS UMi (9/13).
@@ 4.2 Astrometric parameters
The estimated astrometric parameters for ten stars (excluding AR Mon) are listed in Table 4, and the corresponding parallax curves are presented in Appendix C. The position uncertainties of primary calibrators are added to the position uncertainties of targets, so that the uncertainties reflect the errors in absolute position rather than the errors relative to primary calibrators. The uncertainties for astrometric parameters of XY UMa are not reliable values, as discussed in Sect. 3.3, so XY UMa is not included in the discussion below.
@@
The reference epoch is J2023.25. Column “Tech.” denotes calibration technique, PR or sMV. Calibration technique with “” is adopted as the final astrometric result for the star. Correlation coefficient matrices of the adopted results are listed in Appendix D.
In most cases, the sMV provides astrometric parameter uncertainties that are better than, or at least equivalent to, those provided by PR. The performance advantage is also related to the angular separation between the target and the primary calibrator. For stars with a primary calibrator located within (V1355 Ori, FF UMa, RS CVn, RS UMi), the sMV and PR uncertainties are at the same level. While for the other stars, the sMV shows a significant advantage over PR, with two exceptions, HD 8357 and DM UMa.
HD 8357 has a relatively large angular scale of its binary orbit. The semimajor axis projected on the radial direction of its active component km, where is the semimajor axis and is the inclination (1996AJ....112..269F). Calculating with its Gaia parallax mas, the corresponding maximum angular orbital offset would be no less than 0.72 mas, which is non-negligible. Consequently, a better spatial-structure error correction does not substantially improve the final astrometric results.
As mentioned in Sect. 3.3, only two available calibrators are available for DM UMa, and they are both located on the same side of DM UMa in the sky, which leads to a MultiView failure. Therefore, we adopt the PR result as the final astrometric result for DM UMa, whereas the sMV results are used for all other stars.
Overall, a high precision is achieved in most adopted astrometric results. Take parallaxes as an example, the mean ratio is , while the median is , the minimum is , and the maximum is . Proper motions have even better precision levels: from to , of which 9 are better than , benefiting from the long time span (about 3 years). Further improvement requires not only more epochs but also more accurate astrometric models that include binary orbital motions, as discussed above.
@@ 5 Summary
We observed 11 radio stars with the VLBA at C band. With 7 epochs spanning over 3 years, all 11 stars were detected at least twice. The data were calibrated using our pipeline with two techniques: single-calibrator phase referencing and serial MultiView. We successfully derived parallaxes and proper motions for 10 of them. In the best cases, the astrometric parameters reached a precision level of , particularly for proper motions, benefiting from the long time span. The high-quality data produced in this study are an important addition to the available radio star sample for the bright Gaia-CRF validation.
@@ Acknowledgements
The Python pipeline used for calibration can be found at the Astrophysics Source Code Library (2025ascl.soft08016Z), and a “frozen” version used in this paper is available at Zenodo (jingdong_zhang_2025_15030432). The Python program used for astrometric parameter estimation can be found at https://github.com/FrdCHK/mcmc_ppm_estimate, and a “frozen” version used in this paper is available at Zenodo (jingdong_zhang_2025_15031700).
The RFC can be accessed through https://doi.org/10.25966/dhrk-zh08. This work has also made use of the SIMBAD database, operated at CDS, Strasbourg, France (simbad, https://simbad.u-strasbg.fr/simbad/).
@@
This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. The Gaia services (https://gaia.ari.uni-heidelberg.de/index.html) provided by the Astronomisches Rechen-Institut (ARI) of the University of Heidelberg are used in Gaia data retrieval.
@@
Software and Python packages used in this work (in alphabetical order): AIPS (2003ASSL..285..109G), Astropy (astropy2013; astropy2018; astropy2022), Difmap (1997ASPC..125...77S), emcee (2013PASP..125..306F), Matplotlib (2007CSE.....9...90H), Numpy (harris2020array), Pandas (2022zndo...3509134T), ParselTongue (2006ASPC..351..497K), Scipy (2020NatMe..17..261V), and TOPCAT (2005ASPC..347...29T).
@@
The EVN data used in this study can be downloaded from https://archive.jive.nl/scripts/portal.php. The VLBA data used in this study can be downloaded from the data archive of the National Radio Astronomy Observatory (NRAO; https://data.nrao.edu). VLBI images and parallax curves are available at Zenodo (jingdong_zhang_2025_17348486): https://doi.org/10.5281/zenodo.17348486.
Appendix A VLBI Images
Here we put an example (AR Mon) of VLBI images in Fig. 4. Images of all other stars are available at Zenodo (jingdong_zhang_2025_17348486).
Appendix B All measured positions
@@
• All results are produced by the AIPS task JMFIT. Column “Tech.” denotes calibration technique, PR or sMV. denotes peak flux density, and SNR denotes signal-to-noise ratio in image. The coordinates in the first two rows of each star are taken as reference positions for PR and sMV respectively, and the offsets in other rows are relative to them.
Appendix C Estimated parallax curves
@@ Appendix E Self-calibrated images of calibrators
We present self-calibrated images of calibrators that have significant source structures here. Epochs in which all ten VLBA antennas participated are selected for the best UV coverage. Note that each image has a different field of view (FOV). The imaging is done using Difmap (1997ASPC..125...77S).

### v2

@@ 4.1 Gains of MultiView
The performance comparison between the sMV and PR can be conducted using the RMS-based JMFIT error or the fractional flux recovery (FFR) quantity, which is defined as the ratio between the peak brightness in the astrometric image and the self-calibrated image of the same source (Rioja et al., 2017). The radio stars are too weak for self-calibration, so here we use the ratio between the peak flux densities of PR and sMV instead.
@@
Fig. 3 shows the peak flux density ratio of sMV to PR versus the angular separation between the target and the primary calibrator. Similar to the JMFIT error, the sMV reduces flux loss in most cases (sMV/PR > 1, ), and the poor cases are mainly attributed to DM UMa and RS UMi (9/13).
@@ 4.2 Astrometric parameters
The estimated astrometric parameters for ten stars (excluding AR Mon) are listed in Table 4, and the corresponding parallax curves are presented in Appendix C. The position uncertainties of primary calibrators are added to the position uncertainties of targets, so that the uncertainties reflect errors in absolute position rather than offsets to primary calibrators. As discussed in Sect. 3.3, the uncertainties for astrometric parameters of XY UMa are unreliable.
@@
The reference epoch is J2023.25. Column “Tech.” denotes calibration technique, PR or sMV. The calibration technique with “” is adopted as the final astrometric result for each star. Correlation coefficient matrices of the adopted results are listed in Appendix D.
The performance advantage of sMV over PR is related to the angular separation between the target and the primary calibrator. For stars with a primary calibrator located within (V1355 Ori, FF UMa, RS CVn, RS UMi), the sMV and PR uncertainties are at the same level. While for the other stars, the sMV shows a significant advantage over PR, with two exceptions, HD 8357 and DM UMa.
HD 8357 has a relatively large angular scale of its binary orbit. The semimajor axis projected on the radial direction of its active component km, where is the semimajor axis and is the inclination (Fekel, 1996). Calculating with its Gaia parallax mas, the corresponding maximum angular orbital offset would be no less than 0.72 mas, which is non-negligible. Therefore, unless a binary model is adopted, it is difficult to achieve further improvements at the current level of accuracy.
As mentioned in Sect. 3.3, only two available calibrators are available for DM UMa, and they are both located on the same side of DM UMa in the sky, which leads to a MultiView failure. Therefore, we adopt the PR result as the final astrometric result for DM UMa, whereas the sMV results are adopted for all other stars.
Overall, a high precision is achieved in most adopted astrometric results. Take parallaxes as an example, the mean parallax uncertainty is mas, while the median is mas. Parallax is a parameter that is not strongly affected by the systematic bright Gaia-CRF rotation, so a direct comparison between VLBI and Gaia is possible. Fig. 4 shows the comparison between the results and Gaia DR3 parallaxes, for which the zero-point issue has been corrected using the recipe and coefficients provided by Lindegren et al. (2021) and Ding et al. (2024), respectively.
Define the normalized residual between VLBI and Gaia parallaxes as
and its absolute value can be calculated for each star, as shown in Table 5, together with other error statistics. We compared the normalized residuals with and without Gaia parallax zero-point correction, and the correction only reduces the mean absolute value of normalized residual by (, ). Assuming that the VLBI and Gaia measurements are independent, that their uncertainties are reasonable estimates, and that no additional systematic errors exist, should follow a standard Gaussian distribution . Therefore, the probability that is . However, only six of the ten stars have , suggesting that one or both of VLBI and Gaia may have underestimated their uncertainties, or there exist systematic errors.
| Star | |||||
| (mas) | (mas yr-1) | (mas yr-1) | |||
| FF Aqr | 0.072 | 0.043 | 0.061 | 2.144 | 2.212 |
| HD 8357 | 0.366 | 0.223 | 0.483 | 1.782 | 1.638 |
| EI Eri | 0.146 | 0.097 | 0.109 | 3.213 | 3.229 |
| V1859 Ori | 0.085 | 0.056 | 0.034 | 0.803 | 0.240 |
| V1355 Ori | 0.031 | 0.028 | 0.061 | 2.910 | 2.646 |
| XY UMa | 0.571 | 0.427 | 0.328 | 0.011 | 0.078 |
| FF UMa | 0.098 | 0.073 | 0.072 | 1.471 | 1.619 |
| DM UMa | 0.075 | 0.050 | 0.054 | 1.134 | 0.664 |
| RS CVn | 0.104 | 0.051 | 0.107 | 1.300 | 1.099 |
| RS UMi | 0.014 | 0.017 | 0.180 | 3.696 | 2.961 |
| Mean | 0.156 | 0.106 | 0.149 | 1.847 | 1.639 |
| Median | 0.091 | 0.053 | 0.089 | 1.627 | 1.628 |
Columns 2 - 4 list the uncertainties of the adopted parallaxes and proper motions. Column denotes the absolute value of the normalized residual between VLBI and Gaia DR3 parallaxes, while is the normalized residual for Gaia DR3 parallaxes after zero-point correction.
@@ 5 Summary
We observed 11 radio stars with the VLBA at C band. With 7 epochs spanning over 3 years, all 11 stars were detected at least twice. The data were calibrated using our pipeline with two techniques: single-calibrator phase referencing and serial MultiView. We successfully derived parallaxes and proper motions for 10 of them, achieving median uncertainties better than 0.1 mas and 0.1 mas yr-1, respectively. The high-quality data produced in this study are an important addition to the available radio star sample for the bright Gaia-CRF validation, and the frame link results are presented in Zhang et al. (2025a).
@@ Acknowledgements
The Python pipeline used for calibration can be found at the Astrophysics Source Code Library (Zhang et al., 2025c), and a “frozen” version used in this paper is available at Zenodo (Zhang, 2025c). The Python program used for astrometric parameter estimation can be found at https://github.com/FrdCHK/mcmc_ppm_estimate, and a “frozen” version used in this paper is available at Zenodo (Zhang, 2025a).
The RFC can be accessed through https://doi.org/10.25966/dhrk-zh08. This work has also made use of the SIMBAD database, operated at CDS, Strasbourg, France (Wenger et al., 2000, https://simbad.u-strasbg.fr/simbad/).
@@
This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. The Gaia services (https://gaia.ari.uni-heidelberg.de/index.html) provided by the Astronomisches Rechen-Institut (ARI) of the University of Heidelberg are used in Gaia data retrieval. The Python package for Gaia DR3 parallax zero-point correction developed by P. Ramos can be found at https://gitlab.com/icc-ub/public/gaiadr3_zeropoint, and the coefficients used in this paper can be found at https://github.com/yedings/Parallax-bias-correction-in-the-Galactic-plane.
@@
Software and Python packages used in this work (in alphabetical order): AIPS (Greisen, 2003), Astropy (Astropy Collaboration et al., 2013, 2018, 2022), Difmap (Shepherd, 1997), emcee (Foreman-Mackey et al., 2013), Matplotlib (Hunter, 2007), Numpy (Harris et al., 2020), Pandas (The pandas development Team, 2024), ParselTongue (Kettenis et al., 2006), Scipy (Virtanen et al., 2020), and TOPCAT (Taylor, 2005).
@@
The EVN data used in this study can be downloaded from https://archive.jive.nl/scripts/portal.php. The VLBA data used in this study can be downloaded from the data archive of the National Radio Astronomy Observatory (NRAO; https://data.nrao.edu). VLBI images and parallax curves are available at Zenodo (Zhang, 2025b): https://doi.org/10.5281/zenodo.17348486.
Appendix A VLBI images of radio stars
Here we put an example (AR Mon) of VLBI images in Fig. 5. Images of the remaining stars are available at Zenodo (Zhang, 2025b).
Appendix B Measured positions of radio stars
@@
• All results are produced by the AIPS task JMFIT. Column “Tech.” denotes calibration technique, PR or sMV. denotes peak flux density, and SNR denotes signal-to-noise ratio in image. The coordinates in the first two rows of each star are taken as reference positions for PR and sMV, respectively, and the offsets in other rows are relative to them.
Appendix C Estimated parallax curves of radio stars
@@ Appendix D Correlation coefficient matrices of adopted astrometric solutions
The correlation coefficients among the astrometric parameters, (RA, DEC, parallax, and proper motions in the RA and DEC directions), are listed below. In general, the most significant correlations occur between and , and between and , which is physically expected given the coupling between source position and linear proper motion. A relatively high correlation is also present between and the two RA-direction parameters ( and ), because we were primarily sampling at the extremes of parallax sinusoids projected onto the RA direction, as shown in Appendix C. The remaining terms in the correlation coefficient matrices are relatively small.
@@
Since XY UMa was only detected in three epochs, the correlation coefficients between astrometric parameters are extremely high.
@@ Appendix E Self-calibrated images of calibrators
We present self-calibrated images of calibrators that have significant source structures here. Epochs in which all ten VLBA antennas participated are selected for the best UV coverage. Note that each image has a different field of view (FOV). The cleaning and self-calibration are done using Difmap (Shepherd, 1997).